OpenStudy (anonymous):

Find the possible values of b for which the equation 9x^2 + bx + 7 = o has roots which differ by 2.

OpenStudy (anonymous):

This time round i hope someone can post the sum of roots and product of roots method *please*

OpenStudy (badhi):

if the quadratic equation is, $$ax^2+bx+c=0$$and roots are $$\alpha, \; \beta$$ $$\alpha=\frac{-b-\sqrt{\Delta}}{2a},\quad \beta=\frac{-b+\sqrt{\Delta}}{2a}\\ |\alpha-\beta|=\frac{\sqrt{\Delta}}{2a} \quad\text{(for $a>0$)}$$where $$\Delta=b^2-4ac$$

OpenStudy (anonymous):

I have not learnt modulus yet

OpenStudy (anonymous):

going to soon

OpenStudy (badhi):

if you want only from sum and products of the roots, use, $$(\alpha-\beta)^2=(\alpha+\beta)^2-4\alpha\beta$$

OpenStudy (anonymous):

Can you show me that steps please?

OpenStudy (anonymous):

to derive out the possible values of b

OpenStudy (anonymous):

Yours sincerely, very much appreciate it.

OpenStudy (badhi):

you know that $$(\alpha-\beta)^2=2^2=4\\ \alpha+\beta=-\frac ba=-\frac b 9\\ \alpha\beta=\frac ca=\frac 7 9$$ Substituting values to previous given identity, $$4=\left(\frac{-b}{9}\right)^2-4\left(\frac 79 \right)$$ now solve for b. Hope this is helpful

OpenStudy (anonymous):

how did you get (alpha - beta) square ?

OpenStudy (badhi):

In the problem they have given that the difference between the roots is 2. I have taken the roots as alpha and beta Thus we can imply, (alpha - beta)^2 = 4

OpenStudy (anonymous):

Alright, understood then what's the next thing that you did?

OpenStudy (badhi):

As per your information you know how to take the sum and product of the roots from the coefficients of the quadratic equation and simply that's what I've done

OpenStudy (anonymous):

yea that's the first steps of the working.

OpenStudy (anonymous):

Now the third step should be?

OpenStudy (anonymous):

Substituting values to previous given identity, 4=(−b9)2−4(79)

OpenStudy (anonymous):

alright using the (alpha - beta) square = ... formula

OpenStudy (anonymous):

understood.